5x+7y=77
5X+3Y=53

Answer by mananth(16946)

5x+7y=77......................1
5X+3Y=53......................2
..
multiply (1) by -1
we get
-5x-7y=-77
add it to (2)
-5x+5x-7y+3y=-77+53
-4y=-24
/-4
y=6
..
plug value of y in 1
5x+42=77
5x=35
/5
x=7
...
m.ananth@hotmail.ca